# GlobalGeoTree: A Multi-Granular Vision-Language Dataset for Global Tree Species Classification

### v2

@@ 4.1 GeoTreeCLIP model
To establish baseline performance on the GlobalGeoTree dataset, we developed GeoTreeCLIP, a vision-language model specifically designed for tree species classification. The core motivation for adopting the CLIP architecture lies in its ability to learn general and powerful representations through multimodal contrastive learning, which jointly trains on image-text pairs at scale. This approach not only aligns visual and textual modalities but also demonstrates strong transfer capabilities, particularly in zero-shot and few-shot scenarios [14]. Such capabilities are especially valuable given the continuously expanding species catalog and the practical limitations of obtaining exhaustive labeled data for all taxa. These qualities make the CLIP architecture particularly suitable for tree species classification, where leveraging rich textual descriptions (e.g., multi-level taxonomic labels) enhances image understanding and addresses challenges such as rare species representation and complex hierarchical structures. A detailed comparison illustrating the advantages of this contrastive learning approach over a traditional supervised learning paradigm is provided in Appendix D.
@@
  • Visual Encoder: A ViT-B/16 backbone [36] augmented with a temporal attention [37] mechanism to process 12-month Sentinel-2 time series data.
  • Auxiliary Feature Integration: A multi-layer perceptron (MLP) [38, 35] designed to process bioclimatic, soil, and geographic data.
  • Text Encoder: A 77-token causal autoregressive transformer [37, 39] that encodes taxonomic text data across functional type, family, genus, and species levels.
Pretraining details
@@
Both the Visual Encoder and Text Encoder are initialized using OpenAI’s publicly available CLIP checkpoint [13] and further pre-trained on the GlobalGeoTree-6M using Distributed Data Parallel (DDP) across 5 NVIDIA 3090 (24GB) GPUs. We employed a batch size of 384 per GPU, with gradient accumulation over 2 steps, resulting in an effective batch size of 768 per GPU (3840 globally). The AdamW optimizer [40] was used with a base learning rate of $1\times 10^{-5}$ for the visual and auxiliary encoders, and a reduced learning rate of $1\times 10^{-6}$ for the pretrained text encoder. A weight decay of $1\times 10^{-4}$ was applied. A linear learning rate warmup was implemented for the first 5 epochs. Following the warmup, a Cosine Annealing with Warm Restarts [41] scheduler was used, with $T_{0}=10$ epochs and $T_{mult}=2$, and a minimum learning rate of $1\times 10^{-7}$. Gradients were clipped to a maximum L2 norm of 1.0. The loss function was the standard CLIP contrastive loss [42]. Mixed-precision training [43] was enabled. A full 25-epoch training run required approximately 2 days, with peak GPU memory consumption observed at roughly 14 GB per GPU.
@@ 4.2 Experimental setup
We evaluated GeoTreeCLIP against two advanced pretrained vision-language models: the original CLIP [13] and RemoteCLIP [24], a specialized VLM for remote sensing applications. To provide a fair comparison with models not inherently designed for time-series, we adopted an ensemble-like approach for CLIP and RemoteCLIP: features/probabilities were computed for each of the 12 monthly images independently, and the results were then averaged.
@@
Zero-shot evaluation
For zero-shot evaluation, we assessed each model’s ability to classify samples at three taxonomic levels without specific training on the target categories [44].
Few-shot evaluation
For few-shot evaluation, we explored scenarios such as 1-shot learning, where the model is provided with only one labeled example per species. To implement this, we adopted a fine-tuning-based approach [45] using the pre-trained model. Specifically, we randomly sampled $k$ labeled examples per class (e.g., $k=1$, $k=3$) to form the support set and fine-tuned the visual encoder of the pre-trained model on this set. During fine-tuning, most of the visual encoder’s parameters were frozen, with only the last four transformer layers and the classification-related parameters remaining trainable. The text encoder was entirely frozen, leveraging the pre-trained textual embeddings for class labels. The fine-tuning was conducted for 10 epochs to balance adaptation and prevent overfitting due to the small support set. Afterward, the model was evaluated on the query set, which consisted of the remaining examples in the dataset.
@@ 4.3.1 Zero-shot evaluation
The experimental results reveal two key patterns. First, accuracy consistently declines as the taxonomic level becomes finer, reflecting the growing challenge of distinguishing closely related classes. This trend is observed across all models but is particularly pronounced for CLIP and RemoteCLIP, which perform poorly at the genus and species levels. In contrast, GeoTreeCLIP demonstrates stronger performance at these fine-grained levels, likely due to its ability to learn and leverage the hierarchical relationships in taxonomic labels, as supported by feature embedding visualizations (see Appendix E and Figure 7). Second, the significant performance gap between GeoTreeCLIP and the baseline models underscores the importance of domain-specific pretraining. Moreover, its integration of spatiotemporal and multispectral information further enhances its ability as general-purpose models like CLIP and RemoteCLIP struggle to handle. Additional zero-shot benchmark results, including evaluations of SkyCLIP-50 [25] and CLIP-laion-RS, are provided in Appendix F.
@@ 4.3.2 Few-shot evaluation
The poor performance of CLIP and RemoteCLIP can likely be attributed to their design, which is optimized for RGB three-channel data and lacks the capability to process time-series information. Additionally, these models struggle with the small-patch classification tasks required for tree species identification. These limitations further emphasize the importance of introducing this benchmark for the global tree species classification task. More benchmark results on larger evaluation subsets can be found in Table 8 and Table 9 in Appendix C.

### v3

@@ 5.1 GeoTreeCLIP model
To establish a strong baseline on the GlobalGeoTree dataset, we developed GeoTreeCLIP, a vision-language model (VLM) specifically tailored for tree species classification. Our choice of the VLM paradigm was motivated by the inherent challenges of this task: preliminary experiments showed that traditional supervised models struggle with the vast, long-tailed, and hierarchical label space of GlobalGeoTree, especially at the fine-grained genus and species levels (see Sect. 6.1).
In contrast, a VLM framework treats taxonomic labels not as discrete indices, but as structured text. This allows the model to leverage the rich semantic relationships between species, genera, and families. By jointly training on image-text pairs, GeoTreeCLIP learns powerful, transferable representations that align visual features with their corresponding taxonomic context. This approach is particularly effective for zero-shot and few-shot scenarios, which are critical for addressing the immense scale of global biodiversity and the practical impossibility of collecting exhaustive data for all taxa (Stevens et al., 2024).
@@
  • Visual Encoder: A ViT-B/16 backbone (Dosovitskiy et al., 2020) adapted to process the multi-spectral, time-series Sentinel-2 data. To handle the 10-channel, 5x5 pixel input patches, we modify the initial patch embedding layer of the ViT. Each 5x5 patch is first upsampled to 16x16 pixels using bicubic interpolation. The modified embedding layer then processes these 10 channels and projects the entire 16x16 patch into a single visual token. This approach allows the model to capture the holistic spectral information of the patch at each timestep. These 12 monthly tokens are then fed into a temporal attention module (Vaswani et al., 2017) to learn and fuse key phenological patterns across the year. This mechanism allows the model to dynamically weigh the importance of different months, implicitly learning which parts of the phenological cycle are most informative for identifying a particular species.
  • Auxiliary Feature Integration: A multi-layer perceptron (MLP) (Rosenblatt, 1958; Gillespie et al., 2024) designed to process bioclimatic, soil, and geographic data. The MLP consists of several feature-specific linear layers (hidden dimension: 256) followed by LayerNorm and ReLU activation. The encoded features are then fused and projected into a 768-dimensional embedding, which is added to the visual token embedding from the Visual Encoder before the final contrastive projection. This allows the model to integrate environmental context with visual information. We adopt a dual-branch architecture comprising a Visual Encoder and a separate MLP for auxiliary features, rather than stacking all inputs into a single data cube. This design addresses the fundamental heterogeneity of the data modalities. First, Sentinel-2 data consists of a dynamic 12-month time series capturing phenology, whereas auxiliary variables such as bioclimatic and topographic metrics are static. Integrating these static variables into the visual branch would require replicating them across the temporal dimension, introducing significant data redundancy. Second, the spatial resolution of auxiliary variables is typically much coarser ($\sim$1km or 250m) than the 10m Sentinel-2 imagery. Treating them as image bands would necessitate upsampling, creating spatially invariant feature maps that are computationally inefficient for the Vision Transformer. By processing these modalities separately, the MLP branch efficiently encodes environmental context as a conditional prior, which is subsequently fused with spatiotemporal visual features in the shared embedding space.
  • Text Encoder: A 77-token causal autoregressive transformer (Vaswani et al., 2017; Radford et al., 2019) that processes the hierarchical taxonomic labels. For each sample, the text input is structured to convey the full taxonomic lineage, for example, "Evergreen Broadleaf, Family Myrtaceae, Genus Eucalyptus, Species Eucalyptus kybeanensis". By encoding this structured text, the model learns embeddings that capture the relationships between different taxonomic levels. This allows GeoTreeCLIP to understand, for instance, that two different species within the same genus (Eucalyptus) are semantically closer than species from different families. This learned semantic structure is key to its strong generalization capabilities, especially for rare or unseen species.
5.1.1 Pretraining details
@@
Both the Visual Encoder and Text Encoder are initialized using OpenAI’s publicly available CLIP checkpoint (Radford et al., 2021) and further pre-trained on the GlobalGeoTree-6M using Distributed Data Parallel (DDP) across 5 NVIDIA 3090 (24GB) GPUs. We employed a batch size of 384 per GPU, with gradient accumulation over 2 steps, resulting in an effective batch size of 768 per GPU (3840 globally). The AdamW optimizer (Loshchilov and Hutter, 2017) was used with a base learning rate of $1\times 10^{-5}$ for the visual and auxiliary encoders, and a reduced learning rate of $1\times 10^{-6}$ for the pretrained text encoder. A weight decay of $1\times 10^{-4}$ was applied. A linear learning rate warmup was implemented for the first 5 epochs. Following the warmup, a Cosine Annealing with Warm Restarts (Loshchilov and Hutter, 2022) scheduler was used, with $T_{0}=10$ epochs and $T_{mult}=2$, and a minimum learning rate of $1\times 10^{-7}$. Gradients were clipped to a maximum L2 norm of 1.0. The loss function was the standard CLIP contrastive loss (Oord et al., 2018). Mixed-precision training (Micikevicius et al., 2017) was enabled. A full 25-epoch training run required approximately 2 days, with peak GPU memory consumption observed at roughly 14 GB per GPU.
@@ 5.2 Experimental setup
We evaluated GeoTreeCLIP against two advanced pretrained vision-language models: CLIP (Radford et al., 2021) and RemoteCLIP (Liu et al., 2024), a specialized VLM for remote sensing applications. To ensure a fair comparison with models not explicitly designed for time-series data, we employed an ensemble-like approach for CLIP and RemoteCLIP. Specifically, features or probabilities were computed independently for each of the 12 monthly images, and the final results were obtained by averaging these values. Additionally, the results of other baseline models, including SkyCLIP-50 (Wang et al., 2024) and CLIP-laion-RS (He et al., 2024), are provided in Appendix E.
@@
5.2.1 Zero-shot evaluation
For zero-shot evaluation, we assess each model’s ability to classify samples from the evaluation sets (GlobalGeoTree-10kEval, -10kEval-300, and -10kEval-900) without any fine-tuning. Our protocol is framed as an "in-domain" zero-shot classification task, following the setup in recent remote sensing benchmarks (Wang et al., 2024). This approach tests a model’s transfer capability to new, unseen samples of species, rather than to entirely unseen categories. Although the species categories in the evaluation sets are present in the training data, this setup strikes a balance between preserving the model’s ability to classify all tree species after pre-training on GlobalGeoTree-6M and ensuring a valid zero-shot scenario.
5.2.2 Few-shot evaluation
For few-shot evaluation, we explored scenarios such as one-shot learning, where the model is provided with only one labeled example per species. To implement this, we adopted a fine-tuning-based approach (Parnami and Lee, 2022) using the pre-trained model. Specifically, we randomly sampled $k$ labeled examples per class (e.g., $k=1$, $k=3$) to form the support set and fine-tuned the visual encoder of the pre-trained model on this set. During fine-tuning, most of the visual encoder’s parameters were frozen, with only the last four transformer layers and the classification-related parameters remaining trainable. The text encoder was entirely frozen, leveraging the pre-trained textual embeddings for class labels. The fine-tuning was conducted for 10 epochs to balance adaptation and prevent overfitting due to the small support set. Afterward, the model was evaluated on the query set, which consisted of the remaining examples in the dataset.
@@ 5.3.1 Zero-shot evaluation
The experimental results reveal two key patterns. First, accuracy consistently declines as the taxonomic level becomes finer, reflecting the growing challenge of distinguishing closely related classes. This trend is observed across all models but is particularly pronounced for CLIP and RemoteCLIP, which perform poorly at the genus and species levels. In contrast, GeoTreeCLIP demonstrates stronger performance at these fine-grained levels, likely due to its ability to learn and leverage the hierarchical relationships in taxonomic labels, as supported by feature embedding visualizations (see Sect. 6.2 and Figure 8). Second, the significant performance gap between GeoTreeCLIP and the baseline models underscores synergistic effect of domain-specific pretraining and its tailored architecture. Unlike general-purpose models such as CLIP and RemoteCLIP, our approach effectively leverages spatiotemporal and multispectral information to enhance classification capabilities. Additional zero-shot benchmark results, including evaluations of SkyCLIP-50 (Wang et al., 2024) and CLIP-laion-RS (He et al., 2024), are provided in Appendix E.
5.3.2 Performance Analysis by Species Rarity
To assess model robustness against the long-tail distribution inherent in global biodiversity data, we analyzed zero-shot performance at the species level across three frequency groups: Rare, Common, and Frequent. Table 5 details the mean accuracy and standard deviation over 5 runs on the GlobalGeoTree-10kEval benchmark.
Table 5: Zero-shot species-level performance breakdown by rarity group on GlobalGeoTree-10kEval. Results are mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Rarity Group CLIP RemoteCLIP GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Rare 0.00 $\pm$ 0.00 3.61 $\pm$ 0.05 0.00 $\pm$ 0.00 2.43 $\pm$ 0.13 15.25 $\pm$ 0.04 46.65 $\pm$ 0.47
Common 0.64 $\pm$ 0.04 4.52 $\pm$ 0.03 0.00 $\pm$ 0.00 3.45 $\pm$ 0.00 15.93 $\pm$ 0.08 47.51 $\pm$ 0.08
Frequent 3.00 $\pm$ 0.03 9.27 $\pm$ 0.02 3.35 $\pm$ 0.01 12.90 $\pm$ 0.04 17.52 $\pm$ 0.36 51.88 $\pm$ 0.12
The experimental results indicate substantial differences in model generalization across frequency groups. Both baseline models show significant performance degradation on data-scarce classes. Specifically, CLIP yields 0.00% top-1 accuracy for the Rare group, while RemoteCLIP records 0.00% top-1 accuracy across both Rare and Common categories. Conversely, GeoTreeCLIP exhibits consistent performance stability across the frequency spectrum. The model achieves a top-1 accuracy of 15.25% on Rare species, showing a relatively small decrease compared to the 17.52% accuracy observed for Frequent species. This limited performance disparity suggests that GeoTreeCLIP facilitates knowledge transfer through hierarchical taxonomic relationships, enabling the identification of rare species by leveraging features learned from more frequent, related taxa.
5.3.3 Performance Analysis by Geographic Region
To evaluate model generalizability across different geographical regions, we disaggregated the zero-shot results on GlobalGeoTree-10kEval by continent. Table 6 presents the species-level accuracy for each region.
Table 6: Zero-shot species-level performance by continent on GlobalGeoTree-10kEval. Results are mean accuracy (%) $\pm$ standard deviation (%) over 5 runs. (n indicates the sample count per region in the evaluation set).
Region CLIP RemoteCLIP GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Africa ($n=558$) 17.20 $\pm$ 0.05 19.12 $\pm$ 0.05 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 17.31 $\pm$ 0.21 56.63 $\pm$ 1.51
Asia ($n=1120$) 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 4.91 $\pm$ 0.09 8.31 $\pm$ 0.19 39.61 $\pm$ 1.60
Europe ($n=1648$) 0.00 $\pm$ 0.00 19.08 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 19.96 $\pm$ 0.19 49.91 $\pm$ 0.24
North America ($n=4442$) 0.00 $\pm$ 0.00 1.15 $\pm$ 0.02 0.00 $\pm$ 0.00 6.81 $\pm$ 0.01 18.98 $\pm$ 0.33 52.79 $\pm$ 0.45
Oceania ($n=812$) 0.92 $\pm$ 0.06 4.83 $\pm$ 0.07 13.25 $\pm$ 0.00 24.64 $\pm$ 0.18 7.64 $\pm$ 0.11 24.83 $\pm$ 0.02
South America ($n=1150$) 0.00 $\pm$ 0.00 14.76 $\pm$ 0.01 0.00 $\pm$ 0.00 4.60 $\pm$ 0.25 16.83 $\pm$ 0.75 41.89 $\pm$ 1.19
GeoTreeCLIP demonstrates relatively consistent zero-shot capabilities across diverse geographic regions, with top-1 accuracies ranging from 7.64% to 19.96%. The highest performance is observed in Europe and North America, which aligns with the higher density of training samples available for these regions in the GlobalGeoTree dataset. Conversely, performance is lower in Asia and Oceania, likely reflecting the higher species diversity and relative scarcity of labeled data in these areas.
Notably, baseline models exhibit extreme regional bias. CLIP, for example, achieves high accuracy in Africa (17.20%) but drops to 0.00% in most other continents. This suggests that without domain-specific pretraining, models may rely on spurious correlations with broad landscape features (e.g., savannas) rather than learning discriminative species-level traits. GeoTreeCLIP’s ability to maintain performance across all continents highlights its generalization capabilities.
@@ 5.3.4 Few-shot evaluation
The poor performance of CLIP and RemoteCLIP can likely be attributed to their design, which is optimized for RGB three-channel data and lacks the capability to process time-series information. Additionally, these models struggle with the small-patch classification tasks required for tree species identification. These limitations further emphasize the importance of introducing this benchmark for the global tree species classification task. More benchmark results on larger evaluation subsets can be found in Table 15 and Table 16 in Appendix D.
6 In-depth Analysis
6.1 Comparison with Supervised Learning Paradigm
6.1.1 Supervised Model Architecture and Training
6.1.2 Comparison of the Supervised Model
Notably, the GlobalGeoTree-6M dataset was designed to retain nearly all tree species categories, as our goal was to pretrain a model capable of classifying the full spectrum of tree species. Consequently, the zero-shot evaluation here can be seen as measuring the model’s “zero-shot” transfer capability on unseen datasets, aligning with the concept of "in-domain" zero-shot classification defined in Wang et al. (2024).
Table 8: Evaluation on GlobalGeoTree-10kEval. The table compares our vision-language model, GeoTreeCLIP, against the CLIP and a traditional supervised baseline (SupervisedGeoTree). Results are mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. CLIP SupervisedGeoTree GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Family 10.80 $\pm$ 0.03 25.32 $\pm$ 0.05 9.55 $\pm$ 0.10 27.41 $\pm$ 0.07 20.99 $\pm$ 0.28 56.88 $\pm$ 0.42
Genus 1.09 $\pm$ 0.01 9.34 $\pm$ 0.01 1.19 $\pm$ 0.09 8.18 $\pm$ 0.14 18.39 $\pm$ 0.26 50.98 $\pm$ 0.41
Species 1.09 $\pm$ 0.01 7.02 $\pm$ 0.02 0.00 $\pm$ 0.00 0.28 $\pm$ 0.02 16.71 $\pm$ 0.25 47.52 $\pm$ 0.37
In contrast, the proposed GeoTreeCLIP model demonstrates substantial improvements over all other baselines across every taxonomic level. At the family level, GeoTreeCLIP achieves a top-1 accuracy of 20.99%, more than doubling the performance of CLIP and significantly outperforming SupervisedGeoTree. This advantage is even more pronounced at the finer-grained levels: GeoTreeCLIP obtains 18.39% top-1 accuracy for genus and 16.71% for species identification. These results strongly indicate the power of contrastive vision-language learning to align visual features with rich, hierarchical taxonomic text labels. This capability allows GeoTreeCLIP to learn more nuanced and generalizable representations, leading to its superior performance in this challenging zero-shot evaluation. The significant gap, especially compared to SupervisedGeoTree at the species level, highlights the efficacy of the contrastive learning paradigm for handling large, structured label spaces.
6.2 Qualitative Analysis: t-SNE Feature Embeddings Visualization
Figure 8: t-SNE visualization of feature embeddings from GlobalGeoTree-10kEval-300 at different taxonomic levels, comparing GeoTreeCLIP (top row), CLIP (middle row), and SupervisedGeoTree (bottom row). Columns from left to right represent visualizations at the Family level (selected: Anacardiaceae, Berberidaceae, Cactaceae, Hernandiaceae, Rosaceae), Genus level (selected within Rosaceae: Prunus, Pyrus, Rhodotypos, Rubus, Spiraea), and Species level (selected within Prunus: Prunus avium, Prunus caroliniana, Prunus ilicifolia, Prunus laurocerasus, Prunus obtusata).
To qualitatively assess and compare the learned feature representations from different modeling paradigms, we performed t-SNE visualizations (Van der Maaten and Hinton, 2008). We used zero-shot image embeddings (or the final fused visual-auxiliary features for SupervisedGeoTree before the classification heads) extracted from a subset of the GlobalGeoTree-10kEval-300 dataset. This analysis includes our proposed GeoTreeCLIP, the CLIP pretrained by OpenAI, and the SupervisedGeoTree model. Given the large number of classes and the hierarchical nature of the labels, we adopted a selective visualization strategy: first examining embeddings at the family level for five randomly selected families (Anacardiaceae, Berberidaceae, Cactaceae, Hernandiaceae, and Rosaceae); then focusing on five genera within the Rosaceae family (Prunus, Pyrus, Rhodotypos, Rubus, and Spiraea); and finally, visualizing five species within the Prunus genus (Prunus avium, Prunus caroliniana, Prunus ilicifolia, Prunus laurocerasus, and Prunus obtusata).
The comparative t-SNE visualizations are presented in Figure 8. Across all three taxonomic levels (Family, Genus, and Species, shown as columns), GeoTreeCLIP (top row) consistently demonstrates the most effective separation and formation of distinct clusters. At the family level (left column), GeoTreeCLIP clearly distinguishes between the selected families. The CLIP (middle row) exhibits considerable overlap, particularly for the diverse Rosaceae family. SupervisedGeoTree (bottom row) shows some separation but less defined clusters compared to GeoTreeCLIP, with Rosaceae still forming a very broad distribution.
This pattern of superior clustering by GeoTreeCLIP continues at the genus level within Rosaceae (middle column). GeoTreeCLIP forms relatively distinct groups for genera like Prunus, Pyrus, and Rhodotypos. Both CLIP and SupervisedGeoTree struggle more, with CLIP showing a highly condensed and overlapping structure, while SupervisedGeoTree offers some separation but with less clarity than GeoTreeCLIP. The most striking difference is observed at the species level within the Prunus genus (right column). GeoTreeCLIP achieves remarkable separation, forming visually distinct clusters for each of the five Prunus species. In contrast, both CLIP and SupervisedGeoTree largely fail to differentiate these closely related species, with their embeddings heavily intermingled.
6.3 Ablation Study on Input Modalities
To quantify the contribution of different input modalities, we conducted an extensive ablation study. We trained several variants of GeoTreeCLIP (for 10 epochs on GlobalGeoTree-6M) and evaluated their zero-shot performance on the GlobalGeoTree-10kEval benchmark. The results are summarized in Table 9.
Table 9: Ablation study on input modalities. Performance is reported as mean Top-1 accuracy (%) over 5 runs on GlobalGeoTree-10kEval.
Model Variant Family Acc. Genus Acc. Species Acc.
GeoTreeCLIP (Full Input) 18.66% 16.30% 13.95%
- Image Only (10-band S2) 15.83% 14.51% 12.62%
- Auxiliary Only (27 vars) 13.66% 10.50% 9.52%
- RGB + Geolocation Only 7.92% 6.04% 5.50%
- RGB Only 6.38% 5.08% 4.43%
The results lead to several key insights. First, the full input consistently outperforms all ablated versions, highlighting the synergistic effect of integrating spectral-temporal imagery with environmental context. Second, comparing the "Image Only (10-band)" and "RGB Only" variants reveals the critical importance of non-visible bands (e.g., NIR, SWIR) for vegetation analysis, as they provide a substantial performance boost (e.g., 15.83% vs. 6.38% at the family level). Third, adding geolocation to RGB imagery ("RGB + Geolocation") improves performance over "RGB Only", confirming that spatial priors help capture species’ geographical ranges. The reason "Image Only (10-band)" outperforms "RGB + Geolocation" is that the rich information from the 7 additional spectral bands is more discriminative than the signal provided by geolocation alone.